Kc is 4.13x10^-2 for equilibrium 2IBr(g) <> I2(g)+ Br2(g). What are the concentrations of I2(g) and Br2(g) in equilibrium 0.0124 M of IBr(g). Assume that equilibrium is established at the above temperature
By only adding IBr)g)
1
Expert's answer
2020-06-15T14:02:48-0400

c(I2)=c(Br2) = x

 4.13x102 = (x2)/(0,0124)2

x = 0,00252

c(I2)=c(Br2)= 0,00252 M